# Boosting In-Silicon Directed Evolution with Fine-Tuned Protein Language Model and Tree Search

### v2

@@ C.1 Hyperparameters of Fine-Tuning
For the hyperparameters of fine-tuning, we follow the default setting in the fine-tuning script given by the Transformers package (the most widely used Python package for developing large language models). The fine-tuning script using a masked language modeling loss is referred to GitHub. We fine-tune the pretrained protein language models on each task’s homologous sequence dataset for 3 epochs and the batch size per device is set to 8. The learning rate is $5\times 10^{-5}$ and the optimizer is AdamW. The learning rate scheduler type to use is linear while the weight decay is not enabled. These hyperparameter settings in the fine-tuning step are the same for the ESM2-series, ProtBert, and ESM-1b models. These fine-tuning hyperparameters (learning rate, optimizer, epochs, and so on) are default configurations in large language models and work well in our setting. We only set the batch size per device to 8 for low GPU memory cost. The GPU memory should be slightly larger than 32GB to ensure the fine-tuning on the LGK task with the longest protein sequence, with the protein language model size of 650M. We do not fine-tune the ESM2-3B model and the ESM2-15B model, as fine-tuning the two models requires GPUs with much larger memory. At the same time, the results from Table 3 indicate that the small fine-tuned ESM2-35M model owns enough evolution ability to help AlphaDE evolve proteins.
@@ Appendix D AlphaDE with Different EMS2 Model Sizes
Table 3: Comparison with different ESM2-series model sizes. We present the maximum fitness scores obtained in 1000 black-box oracle queries. Results are averaged over five independent trials.
@@ Appendix E Fitness Oracles
The use of TAPE and ESM-1b oracles follows the standard evaluation procedure in previous works (ren_proximal_2022; wang_self-play_2023; qiu_tree_2024). The ESM-1b oracle is used in the avGFP condensing experiment as it supports some special characters, including the token ‘-’. We regard the character token ‘-’ as an amino acid residue deletion and utilize ESM-1b to predict the fitness of the condensed avGFP sequences. These oracle models and their download script could be obtained from GitHub.
@@ Appendix F Ablation Study of AlphaDE
AlphaDE (MCTS with finetuned ESM2-35M model)
@@
Greedy search with finetuned ESM2-35M model
@@
Beam search with finetuned ESM2-35M model
@@ Appendix G Benchmark Experiments with ESM-1b Oracle
We conduct the benchmark experiments with ESM-1b oracle as ren_proximal_2022. Results are shown in Table 6. AlphaDE again beats various competitive baselines in the benchmark setting of EMS-1b oracle.
@@
Task AlphaDE MLDE TreeNeuralTS TreeNeuralUCB EvoPlay PEX
avGFP 3.991 2.443 3.707 3.785 3.835 3.754
UBE2I 3.989 1.434 3.632 3.653 0.647 3.101
@@ H.3 Hyperparameter Sensitivity of Rollout Number
tree depth 1 10 50 100 200 (default) 1000
@@ Appendix I Diversity of Evolved Sequences by AlphaDE
On the other hand, we are also interested in the repetition rate that how often the generated sequences of AlphaDE exist in the task fine-tuning dataset. The repetition rate is calculated by the top $K$ sequences from each trial and there are 5 trials for each task. The top sequences are ranked by the fitness value. Then repetition rate equals to the number of repeated sequences between the top $K$ sequences and the fine-tuning dataset divided by $5\times K$. $K$ is set at 1, 10, 100, and 1000 respectively and the results are shown in Table 11. From Table 11, we see that, in most tasks, the repetition rate maintains at a very low level. This indicates that AlphaDE generates novel sequences instead of repeating the sequences from the fine-tuning dataset. We also note that, as an exception, for the top 1 sequences of task AAV, the repetition rate is relatively high. But this high repetition rate drops as $K$ increases. For example, the repetition rate of the top 10 AAV sequences drops to 26%, which means the high-fitness sequences are mostly different from the fine-tuning sequences.
@@
In the benchmark experiments, almost all the fine-tuned sequence datasets are from deep mutational scanning (DMS), which are not always available for most of the proteins. Therefore, in this section, we introduce the homology searching technique to construct the dataset of homologous sequences for the fine-tuning step in AlphaDE. Next, we use avGFP as an example. Specifically, we use the phmmer homology searching procedure of the biosequence analysis tool HMMER (potter_hmmer_2018) to find homologous sequences of the starting weakest sequence in the avGFP task. We use the default settings of phmmer to search the SwissProt database, UniProt database, Reference Proteomes database, and PDB database. After filtering sequences with the same length as avGFP and removing three duplicate sequences in the DMS dataset, we found 236 unique sequences to construct the avGFP phmmer dataset, which has no overlap with the avGFP DMS dataset. Then we follow the standard AlphaDE fine-tuning step on the avGFP phmmer dataset and perform the MCTS step with the fine-tuned model (here $c$ is set at 1.0 for a better performance). Results are shown in Table 12 below. We see that, AlphaDE, which uses the phmmer homologous sequence dataset for fine-tuning, also achieves superior performance.
@@
Method AlphaDE (DMS) AlphaDE (phmmer) PEX AdaLead TreeNeuralTS
avGFP 3.86 3.83 2.97 2.61 2.44
@@
When initializing the deleted avGFP sequences, we keep the $\beta$-barrel residues and the chromophore related residues. The $\beta$-barrel residues involve the residue 1 to residue 38. The key chromophore residues are residue 65, 66, and 67 (hayes_simulating_2025), and we set the chromophore-related residues to be residue 58 to residue 74. Then we delete half of the left residues, which results in a starting deleted sequence with length 146. In contrast, the wild-type sequence has a length of 238 residues. The wild-type avGFP sequence, the starting deleted sequence, and the final filtered ccGFP1-5 sequences are given in Table 13. When filtering, the amino acid sequences of ccGFP1-5 are fed into AlphaFold 3 server to predict their structures. Then we align these structures with the wild-type structure (PDB ID: 1EMA) by PyMol to calculate the RMSD distances. Although this is a computational proof-of-concept task, it shows the great potential of AlphaDE for different purposes with directed evolution.
@@
At the same time, we also compare with AlphaDE with pretrained protein language model to validate whether it recovers the folded structure of wild-type avGFP. We follow the same pipeline as condensing avGFP with standard AlphaDE. The comparison of the AlphaDE with pretrained ESM2-35M and standard AlphaDE is shown in Table 14 below. It is clear that the condensed avGFPs by AlphaDE with pretrained ESM2-35M have much larger RMSDs, indicating the recovered folded structures are less similar to the wild-type avGFP structure than the standard AlphaDE.
Table 14: Comparison of AlphaDE with pretrained ESM2-35M and standard AlphaDE with fine-tuned ESM2-35M for condensing avGFP. The metric is RMSD to the wild-type avGFP (PDB ID: 1EMA).
@@
Here we give the running time of the fine-tuning step and the MCTS inference step. The protein language model here is ESM2-35M, which is the default configuration in AlphaDE. The running time is shown in Table 15. For the fine-tuning step, the running time mainly depends on the number of sequences in the dataset. We use three NVIDIA GPUs for this fine-tuning step and one NVIDIA GPU for the MCTS inference step. When fine-tuning, the GPU memory cost depends on the size of protein language models and the length of protein amino acid sequences.
@@
Meanwhile, we additionally provide the computational cost of competitive baselines such as TreeNeuralTS and TreeNeuralUCB for a comparison in the task avGFP. We also compare with EvoPlay. The results are included in Table 16. As AlphaDE utilizes the protein language model, it is expected to take a longer time to run. We also see that, the running time of AlphaDE is acceptable, compared with other baselines.
Table 16: Running time of AlphaDE and different baselines in the task of avGFP. The unit of running time is the hour. The resulting running hour values are averaged over 5 trials.

### v3

@@ C.1 Hyperparameters of Fine-Tuning
For the hyperparameters of fine-tuning, we follow the default setting in the fine-tuning script given by the Transformers package (the most widely used Python package for developing large language models). The fine-tuning script using a masked language modeling loss is available on GitHub. We fine-tune the pretrained protein language models on each task’s homologous sequence dataset for 3 epochs and the batch size per device is set to 8. The learning rate is $5\times 10^{-5}$ and the optimizer is AdamW. The learning rate scheduler type to use is linear while the weight decay is not enabled. These hyperparameter settings in the fine-tuning step are the same for the ESM2-series, ProtBert, and ESM-1b models. These fine-tuning hyperparameters (learning rate, optimizer, epochs, and so on) are default configurations in large language models and work well in our setting. We only set the batch size per device to 8 for low GPU memory cost. The GPU memory should be slightly larger than 32GB to ensure the fine-tuning on the LGK task with the longest protein sequence, with the protein language model size of 650M. We do not fine-tune the ESM2-3B model and the ESM2-15B model, as fine-tuning the two models requires GPUs with much larger memory. At the same time, the results from Table 3 indicate that the small fine-tuned ESM2-35M model owns enough evolution ability to help AlphaDE evolve proteins.
@@ Appendix D AlphaDE with Different EMS2 Model Sizes
Table 3: Performance of AlphaDE with different ESM2 model sizes. We present the maximum fitness scores obtained in 1000 black-box oracle queries. Results are averaged over five independent trials.
@@ Appendix E Fitness Oracles
The use of TAPE and ESM-1b oracles follows the standard evaluation procedure in previous works (ren_proximal_2022; wang_self-play_2023; qiu_tree_2024). The ESM-1b oracle is used in the avGFP condensing experiment as it supports some special characters, including the token ‘-’. We regard the character token ‘-’ as an amino acid residue deletion and utilize ESM-1b to predict the fitness of the condensed avGFP sequences. These oracle models and their download script can be obtained from GitHub.
@@ Appendix F Ablation Study of AlphaDE
AlphaDE (MCTS with fine-tuned ESM2-35M model)
@@
Greedy search with fine-tuned ESM2-35M model
@@
Beam search with fine-tuned ESM2-35M model
@@ Appendix G Benchmark Experiments with ESM-1b Oracle
We conduct the benchmark experiments with ESM-1b oracle taken from ren_proximal_2022. Results are shown in Table 6. AlphaDE again beats various competitive baselines in the benchmark setting of EMS-1b oracle.
@@
Task AlphaDE LatentDE MLDE TreeNeuralTS TreeNeuralUCB EvoPlay PEX
avGFP 3.991 3.677 2.443 3.707 3.785 3.835 3.754
UBE2I 3.989 2.360 1.434 3.632 3.653 0.647 3.101
@@ H.3 Hyperparameter Sensitivity of Rollout Number
rollout number 1 10 50 100 200 (default) 1000
@@ Appendix I Diversity of Evolved Sequences by AlphaDE
On the other hand, we are also interested in the repetition rate that how often the generated sequences of AlphaDE exist in each task’s fine-tuning dataset. The repetition rate is calculated by the top $K$ sequences from each trial, and there are 5 trials for each task. The top sequences are ranked by the fitness value. Then the repetition rate equals to the number of repeated sequences between the top $K$ sequences and the fine-tuning dataset divided by $5\times K$. $K$ is set at 1, 10, 100, and 1000 respectively, and the results are shown in Table 11. From Table 11, we see that, in most tasks, the repetition rate maintains at a very low level. This indicates that AlphaDE generates novel sequences instead of repeating the sequences from the fine-tuning dataset. We also note that, as an exception, for the top 1 sequences of task AAV, the repetition rate is relatively high. But this high repetition rate drops as $K$ increases. For example, the repetition rate of the top 10 AAV sequences drops to 26%, which means the high-fitness sequences are mostly different from the fine-tuning sequences.
@@
Method AlphaDE (DMS) AlphaDE (phmmer) PEX (CNN) AdaLead TreeNeuralTS
avGFP 3.86 3.83 3.30 1.81 2.44
In the benchmark experiments, almost all the fine-tuned sequence datasets are from deep mutational scanning (DMS), which are not always available for most proteins. Therefore, in this section, we introduce the homology searching technique to construct the dataset of homologous sequences for the fine-tuning step in AlphaDE. Next, we use avGFP as an example. Specifically, we use the phmmer homology searching procedure of the biosequence analysis tool HMMER (potter_hmmer_2018) to find homologous sequences of the starting weakest sequence in the avGFP task. We use the default settings of phmmer to search the SwissProt database, UniProt database, Reference Proteomes database, and PDB database. After filtering sequences with the same length as avGFP and removing three duplicate sequences in the DMS dataset, we found 236 unique sequences to construct the avGFP phmmer dataset, which has no overlap with the avGFP DMS dataset. Then we follow the standard AlphaDE fine-tuning step on the avGFP phmmer dataset and perform the MCTS step with the fine-tuned model (here $c$ is set at 1.0 for a better performance). Results are shown in Table 12 below. We see that, AlphaDE, which uses the phmmer homologous sequence dataset for fine-tuning, also achieves superior performance.
@@
When initializing the deleted avGFP sequences, we keep the $\beta$-barrel residues and the chromophore-related residues. The $\beta$-barrel residues involve the residue 1 to residue 38. The key chromophore residues are residue 65, 66, and 67 (hayes_simulating_2025), and we set the chromophore-related residues to be residue 58 to residue 74. Then we delete half of the remaining residues, which results in a starting deleted sequence with length 146. In contrast, the wild-type sequence has a length of 238 residues. The wild-type avGFP sequence, the starting deleted sequence, and the final filtered ccGFP1-5 sequences are given in Table 13. When filtering, the amino acid sequences of ccGFP1-5 are fed into AlphaFold 3 server to predict their structures. Then we align these structures with the wild-type structure (PDB ID: 1EMA) by PyMol to calculate the RMSD distances. Although this is a computational proof-of-concept task, it shows the great potential of AlphaDE for different purposes with directed evolution.
@@
At the same time, we also compare with AlphaDE with the pretrained protein language model to validate whether it recovers the folded structure of wild-type avGFP. We follow the same pipeline as condensing avGFP with standard AlphaDE. The comparison of the AlphaDE with pretrained ESM2-35M and standard AlphaDE is shown in Table 14 below. It is clear that the condensed avGFPs by AlphaDE with pretrained ESM2-35M have much larger RMSDs, indicating the recovered folded structures are less similar to the wild-type avGFP structure than the standard AlphaDE.
Table 14: Comparison of AlphaDE with pretrained ESM2-35M and with fine-tuned ESM2-35M for condensing avGFP. The metric is RMSD to the wild-type avGFP (PDB ID: 1EMA).
@@
Here, we give the running time of the fine-tuning step and the MCTS inference step. The protein language model here is ESM2-35M, which is the default configuration in AlphaDE. The running time is shown in Table 15. For the fine-tuning step, the running time mainly depends on the number of sequences in the dataset. We use three NVIDIA GPUs for this fine-tuning step and one NVIDIA GPU for the MCTS inference step. When fine-tuning, the GPU memory cost depends on the size of protein language models and the length of protein amino acid sequences.
@@
Meanwhile, we additionally provide the computational cost of competitive baselines such as TreeNeuralTS and TreeNeuralUCB for a comparison in the task avGFP. We also compare with EvoPlay. The results are included in Table 16. As AlphaDE utilizes the protein language model, it is expected to take a longer time to run. We also see that the running time of AlphaDE is acceptable, compared with other baselines.
Table 16: Running time of AlphaDE and different baselines in avGFP averaged over 5 trials. The unit is the hour.